# A Monocular Wide-FoV SLAM System Based on Point Uncertainty and Pose Uncertainty

### v1

@@ Abstract
This paper presents P2U-SLAM, a visual Simultaneous Localization And Mapping (SLAM) system with a wide Field of View (FoV) camera, which utilizes pose uncertainty and point uncertainty. While the wide FoV enables considerable repetitive observations of historical map points for matching cross-view features, the data properties of the historical map points and the poses of historical keyframes have changed during the optimization process. The neglect of data property changes triggers the absence of a partial information matrix in optimization and leads to the risk of long-term positioning performance degradation. The purpose of our research is to reduce the risk of the wide field of view visual input to the SLAM system. Based on the conditional probability model, this work reveals the definite impact of the above data properties changes on the optimization process, concretizes it as point uncertainty and pose uncertainty, and gives a specific mathematical form. P2U-SLAM respectively embeds point uncertainty and pose uncertainty into the tracking module and local mapping, and updates these uncertainties after each optimization operation including local mapping, map merging, and loop closing. We present an exhaustive evaluation in $27$ sequences from two popular public datasets with wide-FoV visual input. P2U-SLAM shows excellent performance compared with other state-of-the-art methods. The source code will be made publicly available at https://github.com/BambValley/P2U-SLAM.
Index Terms:
Wide-FoV cameras, uncertainty, pose optimization, visual SLAM.
@@ I Introduction
Intensive investigation on Visual Simultaneous Localization and Mapping (SLAM), a method utilizing sensors to estimate agent pose and establish a map of environment [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], is emerging for decades, in the field of autonomous navigation and intelligent transportation system. Given the robustness of many research works on SLAM suffering from rapid motion and dynamic scenarios, wide Field of View (FoV) cameras (Fig. 1), such as panoramic annular cameras or fisheye cameras, have become a kind of important sensor for SLAM application in light of their natural wider overlap area across frames [15, 16, 17, 18, 19, 20, 21, 22].
@@
There have been multiple visual SLAM (vSLAM) frameworks [23, 19, 24, 20, 21, 22, 25, 26] owning the ability to apply wide-FoV cameras and obtain progress of robustness for enhancing pose estimation in real-world transportation scenarios. These frameworks have realized the adaptation of the nonlinear wide-FoV camera models and the corresponding epipolar constraint due to nonlinear models.
@@
Figure 1: An example of the wide-FoV lens used in transportation systems. a) The aerial platform used for data collection in the PALVIO dataset [26], equipped with two wide-FoV Panoramic Annular Lens (PAL) cameras; b) The output trajectories of several SLAM/VIO algorithms on the ID01 sequence of the PALVIO dataset, where the blue and green trajectories are the results of the P2U-SLAM proposed in this paper. In addition, (w.) indicates running with the support of point uncertainty and pose uncertainty and (w./o.) indicates running without that support.
@@
However, the impact of the switch from ordinary pinhole cameras to wide-FoV cameras, which involves many aspects of SLAM, should be more comprehensively investigated. In the process of matching features in the front-end of vSLAM, different FoV distribution of the same feature induces a huge difference of distortion in co-visible frames [27, 28]. This difference of distortion results in a change of the neighborhood grayscale arrangement of the 2D feature in a wide-FoV image [29], and leads to weak visual constraints for feature-based methods. On the other hand, a wider overlap area produces more observations of 3D points between keyframes, from which Bundle Adjustment (BA) in the SLAM system could benefit [20]. At the same time, more weak constraints might be introduced as the same 3D point could be observed by two keyframes that are much farther apart from each other, and the uncertainty of 3D points and co-visible keyframe poses will be more crucial when the camera is re-approaching a historical observation point [30]. Though those weak constraints are enhanced or eliminated within several turns of BA optimization, the accuracy and efficiency of pose estimation have been influenced [31]. Some frameworks [32, 33, 34, 2, 20] of direct or semi-direct visual SLAM suppress the impact of point uncertainty through a Bayesian framework before these immature points go to the process of BA, yet, the Bayesian framework is only effective within a short-term window of a few seconds. It remains powerless against the effects of point re-observation in medium or long-term tracking, which could last for hours.
@@
In view of this, we propose P2U-SLAM, a feature-based wide-FoV SLAM framework based on point uncertainty and pose uncertainty. Unlike the fixed round 2D covariance used for conventional SLAM frameworks such as ORB-SLAM3 [23], PAL-SLAM [22], and so on, P2U-SLAM utilizes the ellipsoidal 3D point uncertainty, which can be combined with the Jacobian matrix of camera model to adapt to changes in neighborhood distribution caused by distortion of wide-FoV images. In the meantime, point uncertainty and pose uncertainty can acquire historical observation errors to instruct the direction of optimization. They are quantified as a covariance matrix and integrated into the information matrix of BA so that weak constraints introduced by weak feature correspondence from a wide overlap area will be suppressed at any time of the BA process. P2U-SLAM adopts a multi-threading solution to ensure operation efficiency, mainly including tracking, local mapping, and loop detection threads. Point uncertainty and pose uncertainty are applied in the tracking thread and local mapping thread respectively to improve the accuracy and efficiency of the optimization process.
@@
    The calculation approach of point/pose uncertainty is proposed to overcome the shortcoming that the point/pose uncertainty introduces weak constraints for a wider overlap area, which is quantified as a 3D/6D covariance matrix, and it can flexibly converse to the 2D version with the Jacobian matrix of camera model’s project function.
@@
The remainder of this paper is organized as follows. Sec. II provides a brief review of related works. Sect. III introduces the wide-FoV camera models P2U-SLAM supports. Sect. IV explains why point uncertainty and pose uncertainty are important and how to use them in detail. A system overview of P2U-SLAM is provided in Sec. V. In Sec. VI, a comprehensive set of experiments are performed to evaluate our algorithm. Finally, the conclusions are drawn and future research directions are pointed out in Sec. VII.
@@ II-A VO/VIO and SLAM
Visual odometry is the process of computing the six Degree-of-Freedom (DoF) pose by using vision measurement units. Engel et al. [4] proposed Direct Sparse Odometry (DSO), a classic VO framework using direct methods, which directly exploits pixel intensity values instead of artificial vision features. Forster et al. [2] proposed Semi-direct monocular Visual Odometry (SVO), which realized a combination of direct and feature-based methods. Visual-inertial odometry is a VO framework that utilizes vision sensors and inertial measurement units. Qin et al. proposed VINS-Mono [9], a robust and universal monocular vision inertial state estimator. Then, SVO2.0 [3], the visual-inertial version of SVO, has been designed for monocular and multi-camera systems. SLAM can be regarded as the VO/VIO with loop closure and global optimization after years of development. Mur-Artal et al. designed a novel monocular SLAM system named ORB-SLAM [7]. Nowadays, ORB-SLAM has been upgraded to ORB-SLAM3 [23], an accurate open-source library for vision, visual-inertial, and multi-map SLAM. Similarly, VINS-Fusion [35, 36] has realized the ability of loop closure to become a complete SLAM system based on VINS-Mono.
Nevertheless, most of the above frameworks are not able to work well with wide-FoV camera systems owning $360^{\circ}$ omnidirectional FoV in actual deployment in autonomous navigation systems. Therefore, there are plenty of recent works aiming at the adaptation of nonlinear wide-FoV cameras, which even have a negative plane with FoV beyond $180$°.
@@ II-B Wide-FoV SLAM
The wide-FoV SLAM exhibits high stability of localization in situations such as rapid motion and scene transitions. Common wide-FoV lenses mainly include fisheye lens [37], catadioptric cameras [38], multi-camera panoramic imaging systems, and Panoramic Annular Lens (PAL) cameras [39, 40]. The Omnidirectional DSO [24] was designed by Matsuki et al., and it combines the MEI camera model [41] and DSO to build a direct VO that supports the omnidirectional camera. ORB-SLAM3 [23] utilized the Kannala-Brandt fisheye model [42] based on their previous work to support fisheye lens. Panoramic Visual Odometry (PVO) [43] was proposed for the Ricoh Theta V panoramic camera with a panoptic semantic map generation. OpenVSLAM [44] was proposed by Sumikura et al., which can adapt to panoramic cameras based on ORB-SLAM2 [8]. ROVINS [45] used four fisheye cameras and Inertial Measurement Units (IMUs) to achieve an omnidirectional multi-view VIO, which is robust in fast motion or sudden illumination changes.
@@
Due to the low resolution of the fisheye lens at the edge of the FoV, many classic SLAM frameworks do not support the FoV from the negative half-plane beyond $180$°, although they support fisheye camera models. However, PAL tends to maintain low radial distortion in the FoV from the negative half plane, which is comparable to the central FoV, as illustrated in Fig. 2. Therefore, the mentioned classic SLAM frameworks above cannot adapt to the ultra-large FoV of PAL. Chen et al. [20] proposed PALVO, introducing the Taylor camera model to VO methods and realizing high robustness in rapid motion and dynamic scenarios. PALVO is then updated to PA-SLAM [21] with extra ORB features [46, 47] extraction and the help of DBoW2 [48] to incorporate loop-closure module and become a SLAM system. Hu et al. [19] proposed PAL-DSO, which extends the direct methods to PAL cameras. Wang et al. [22] proposed PAL-SLAM to establish a complete visual SLAM framework based on PAL cameras and feature-based methods. Recently, Wang et al. updated their work to PAL-SLAM2 [25], implementing support for IMU input. LF-VIO [26] proposed a real-time VIO framework to compare the impact of visual input from different FoVs on the VIO system.
@@
Wide-FoV input has been proven to be crucial for improving robustness and accuracy in positioning systems [49, 50, 51]. Many extension works based on wide-FoV SLAM and visual odometry have appeared recently. Wang et al. introduced loop closure modules and geodesic constraints to improve the positioning ability atop LF-VIO [52, 53]. Xie et al.implemented a lightweight panoramic depth completion network to acquire dense depth from a back-to-back dual fisheye camera setup and generate a dense 3D map [54]. Xiao et al. made use of panoramic semantic information to overcome the issue of inconsistent visual contents from different viewpoints in loop closure [55].
@@
Although there have been plenty of VIO and SLAM frameworks that support wide-FoV cameras, almost all of them pay no attention to the uncertainty of variables observed multiple times by wide-FoV cameras and the corresponding variables’ properties changes during the location process. Some methods [9, 19, 26, 24] cannot update the uncertainty of variables continuously due to the limitation of direct methods or the optical flow tracking in the matching process. We point out that the methods based on descriptors have the potential to calculate and utilize the uncertainty of 3D points and 6D poses simultaneously. To make use of the potential of descriptors and overcome the limitations introduced by changes in variables’ properties, we propose the first visual wide-FoV SLAM system utilizing point- and pose uncertainty based on ORB features.
@@ III Camera Model
The camera model describes the measurement process of the 3D coordinate of a map point and outputs a projection pixel on the image plane as its measurement result. Many SLAM systems rectify the images to adapt to the pinhole camera model, while the way of distortion rectification cannot be flexibly transferred to a wide-FoV camera, whose FoV can even surpass $180$°. Equidistance projection and other non-perspective projection camera models are more suitable for calibration of wide-FoV camera [42] since the rectification of wide-FoV images is unnecessary. Considering that the wide-FoV camera intrinsic parameters provided by different datasets may come from different camera models, we provide two popular camera models for wide-FoV cameras, which are the Taylor camera model [56] and the Kannala-Brandt fisheye model [42].
@@ III-A Taylor Camera Model
As described in [19, 56], the projection function between a map point $\mathbf{p}$ with camera coordinate $(x,y,z)^{T}$ and its projection pixel $\mathbf{u}_{\pi}$ with coordinate $(u,v)^{T}$ can be written as
@@
$\rho(\theta)=a_{0}+a_{1}\cdot\theta^{1}+a_{2}\cdot\theta^{2}+a_{3}\cdot\theta^{3}+...,$
@@
$\mathbf{\overline{p}}=\pi^{-1}(\mathbf{u}_{\pi})=\frac{\mathbf{p}^{\prime}}{||\mathbf{p}^{\prime}||},$ (3)
@@
$z_{p}(r)=b_{0}+b_{1}\cdot r^{1}+b_{2}\cdot\ r^{2}+b_{3}\cdot r^{3}+...,$
@@
It should be noted that the Jacobian matrix of the camera projection function can be calculated with chain rules as Eq. (5), which matters in motion estimation and calculation of point uncertainty.

### v2

@@ Abstract
This paper presents P2U-SLAM, a visual Simultaneous Localization And Mapping (SLAM) system with a wide Field of View (FoV) camera, which utilizes pose uncertainty and point uncertainty. While the wide FoV enables considerable repetitive observations of historical map points for matching cross-view features, the data properties of the historical map points and the poses of historical keyframes have changed during the optimization process. The neglect of data property changes results in the lack of partial information matrices in optimization, increasing the risk of long-term positioning performance degradation. The purpose of our research is to mitigate the risks posed by wide-FoV visual input to the SLAM system. Based on the conditional probability model, this work reveals the definite impacts of the above data properties changes on the optimization process, concretizes these impacts as point uncertainty and pose uncertainty, and gives their specific mathematical form. P2U-SLAM embeds point uncertainty into the tracking module and pose uncertainty into the local mapping module respectively, and updates these uncertainties after each optimization operation including local mapping, map merging, and loop closing. We present an exhaustive evaluation on $27$ sequences from two popular public datasets with wide-FoV visual input. P2U-SLAM shows excellent performance compared with other state-of-the-art methods. The source code will be made publicly available at https://github.com/BambValley/P2U-SLAM.
@@ I Introduction
Intensive investigation on Visual Simultaneous Localization and Mapping (SLAM), a method utilizing sensors to estimate agent pose and establish a map of environment [cadena2016past, forster2014svo, forster2016svo, engel2017direct, 2017Stereo, zubizarreta2020direct, mur2015orb, mur2017orb, qin2018vins, 10461980, 9994237, 10591477, 10542573, 10260268, Ge2024PIPOSLAMLV, Xu2022D2SLAMDA, Boche2025OKVIS2XOK], is emerging for decades, in the field of autonomous navigation and intelligent transportation system. Given the robustness of many research works on SLAM suffering from rapid motion and dynamic scenarios, wide Field of View (FoV) cameras (Fig. 1), such as panoramic annular cameras or fisheye cameras, have become an important type of sensor for SLAM application in light of their natural wider overlap area across frames [7009773, 9554752, 8813985, 10610351, hu2019indoor, chen2019palvo, chen2021panoramic, wang2022pal].
@@
There have been multiple visual SLAM (vSLAM) frameworks [campos2021orb, hu2019indoor, matsuki2018omnidirectional, chen2019palvo, chen2021panoramic, wang2022pal, wang2024pal, wang2022lf] owning the ability to apply wide-FoV cameras and obtain progress of robustness for enhancing pose estimation in real-world transportation scenarios. These frameworks have realized the adaptation of the nonlinear wide-FoV camera models and the corresponding epipolar constraint due to nonlinear models.
@@
Figure 1: An example of the wide-FoV lens used in transportation systems. a) The aerial platform used for data collection in the PALVIO dataset [wang2022lf], equipped with two wide-FoV Panoramic Annular Lens (PAL) cameras; b) The output trajectories of several SLAM/VIO algorithms on the ID01 sequence of the PALVIO dataset, where the blue and green trajectories are the results of the P2U-SLAM proposed in this paper. In addition, (w.) indicates running with the support of point uncertainty and pose uncertainty, and (w./o.) indicates running without that support.
@@
However, the impact of the switch from ordinary pinhole cameras to wide-FoV cameras, which involves many aspects of SLAM, should be more comprehensively investigated. In the process of matching features in the front-end of vSLAM, different FoV distribution of the same feature induces a huge difference of distortion in co-visible frames [zhao2015sphorb, urban2017mdbrief]. This difference of distortion results in a change of the neighborhood grayscale arrangement of the 2D feature in a wide-FoV image [zhu2018robust], and leads to weak visual constraints for feature-based methods. On the other hand, a wider overlap area produces more observations of 3D points between keyframes, from which Bundle Adjustment (BA) in the SLAM system could benefit [chen2019palvo]. At the same time, additional weak constraints might be introduced as the same 3D point could be observed by two keyframes that are much farther apart from each other, and the uncertainty of 3D points and co-visible keyframe poses will be more crucial when the camera is re-approaching a historical observation point [dai2020rgb]. Though those weak constraints are enhanced or eliminated within several turns of BA optimization, the accuracy and efficiency of pose estimation have been influenced [park2022keeping]. Some frameworks [2003Real, 2007MonoSLAM, 2007A, forster2014svo, chen2019palvo] of direct or semi-direct visual SLAM suppress the impact of point uncertainty through a Bayesian framework before these immature points are fed into the process of BA, yet, the Bayesian framework is only effective within a short-term window of a few seconds. It remains powerless against the effects of point re-observation in medium- or long-term tracking, which could last for hours.
@@
In view of this, we propose P2U-SLAM, a feature-based wide-FoV SLAM framework based on point uncertainty and pose uncertainty. Unlike the fixed round 2D covariance used for conventional SLAM frameworks such as ORB-SLAM3 [campos2021orb], PAL-SLAM [wang2022pal], and so on, P2U-SLAM utilizes the ellipsoidal 3D point uncertainty, which can be combined with the Jacobian matrix of the camera model to adapt to changes in neighborhood distribution caused by distortion of wide-FoV images. In the meantime, point uncertainty and pose uncertainty can acquire historical observation errors to instruct the direction of optimization. They are quantified as a covariance matrix and integrated into the information matrix of BA so that weak constraints introduced by weak feature correspondence from a wide overlap area will be suppressed at any time of the BA process. P2U-SLAM adopts a multi-threading solution to ensure operation efficiency, mainly including tracking, local mapping, and loop detection threads. Point uncertainty and pose uncertainty are applied in the tracking thread and local mapping thread respectively to improve the accuracy and efficiency of the optimization process.
@@
    The calculation approach of point/pose uncertainty is proposed to overcome the shortcoming that the point/pose uncertainty introduces weak constraints for a wider overlap area, which is quantified as a 3D/6D covariance matrix, and it can flexibly converse to the 2D version with the Jacobian matrix of the camera model’s project function.
@@
The remainder of this paper is organized as follows. Sec. II provides a brief review of related works. SecIII introduces the wide-FoV camera models P2U-SLAM supports. SecIV explains why point uncertainty and pose uncertainty are important and how to use them in detail. A system overview of P2U-SLAM is provided in Sec. V. In Sec. VI, a comprehensive set of experiments is performed to evaluate our algorithm. Finally, the conclusions are drawn and future research directions are pointed out in Sec. VII.
@@ II-A VO/VIO and SLAM
Visual odometry is the process of computing the six Degree-of-Freedom (DoF) pose by using vision measurement units. Engel et al. [engel2017direct] proposed Direct Sparse Odometry (DSO), a classic VO framework using direct methods, which directly exploits pixel intensity values instead of artificial vision features. Forster et al. [forster2014svo] proposed Semi-direct monocular Visual Odometry (SVO), which realized a combination of direct and feature-based methods. Visual-inertial odometry is a VO framework that utilizes vision sensors and inertial measurement units. Qin et al. proposed VINS-Mono [qin2018vins], a robust and universal monocular vision inertial state estimator. Then, SVO2.0 [forster2016svo], the visual-inertial version of SVO, has been designed for monocular and multi-camera systems. SLAM can be regarded as the VO/VIO with loop closure and global optimization after years of development. Mur-Artal et al. designed a novel monocular SLAM system named ORB-SLAM [mur2015orb]. Nowadays, ORB-SLAM has been upgraded to ORB-SLAM3 [campos2021orb], an accurate open-source library for vision, visual-inertial, and multi-map SLAM. Similarly, VINS-Fusion [qin2018online, qin2019vins] has realized the ability of loop closure to become a complete SLAM system based on VINS-Mono.
Nevertheless, most of the above frameworks are not able to work well with wide-FoV camera systems featuring $360^{\circ}$ omnidirectional FoV in actual deployment in autonomous navigation systems. Therefore, there are plenty of recent works aiming at the adaptation of nonlinear wide-FoV cameras, which even have a negative half plane with FoV beyond 180180°.
@@ II-B Wide-FoV SLAM
The wide-FoV SLAM exhibits high stability of localization in situations such as rapid motion and scene transitions. Common wide-FoV lenses mainly include fisheye lenses [9913937], catadioptric cameras [1460954], multi-camera panoramic imaging systems, and Panoramic Annular Lens (PAL) cameras [9927463, Wang2022HighperformancePA]. The Omnidirectional DSO [matsuki2018omnidirectional] was designed by Matsuki et al., and it combines the MEI camera model [geyer2000unifying] and DSO to build a direct VO that supports the omnidirectional camera. ORB-SLAM3 [campos2021orb] utilized the Kannala-Brandt fisheye model [kannala2006generic] based on their previous work to support the fisheye lens. Panoramic Visual Odometry (PVO) [8610700] was proposed for the Ricoh Theta V panoramic camera with a panoptic semantic map generation. OpenVSLAM [2019arXiv191001122S] was proposed by Sumikura et al., which can adapt to panoramic cameras based on ORB-SLAM2 [mur2017orb]. ROVINS [9144432] used four fisheye cameras and Inertial Measurement Units (IMUs) to achieve an omnidirectional multi-view VIO, which is robust in fast motion or sudden illumination changes.
@@
Due to the low resolution of the fisheye lens at the edge of the FoV, many classic SLAM frameworks do not support the FoV from the negative half-plane beyond 180180°, although they support fisheye camera models. However, PAL tends to maintain low radial distortion in the FoV from the negative half plane, which is comparable to the central FoV, as illustrated in Fig. 2. Therefore, the mentioned classic SLAM frameworks above cannot adapt to the ultra-large FoV of PAL. Chen et al. [chen2019palvo] proposed PALVO, introducing the Taylor camera model to VO methods and realizing high robustness in rapid motion and dynamic scenarios. PALVO is then updated to PA-SLAM [chen2021panoramic] with extra ORB features [2010BRIEF, 2011ORB] extraction and the help of DBoW2 [galvez2012bags] to incorporate a loop-closure module and become a SLAM system. Hu et al. [hu2019indoor] proposed PAL-DSO, which extends the direct methods to PAL cameras. Wang et al. [wang2022pal] proposed PAL-SLAM to establish a complete visual SLAM framework based on PAL cameras and feature-based methods. Recently, Wang et al. updated their work to PAL-SLAM2 [wang2024pal], implementing support for IMU input.
@@
Wide-FoV input has been proven to be crucial for improving robustness and accuracy in positioning systems [wu2023_360vio, ahmadi2023hdpv, yang2024mcov]. Many extension works based on wide-FoV SLAM and visual odometry have appeared recently. Xie et al. implemented a lightweight panoramic depth completion network to acquire dense depth from a back-to-back dual fisheye camera setup and generate a dense 3D map [xie2024omnidirectional_dense]. Xiao et al. made use of panoramic semantic information to overcome the issue of inconsistent visual contents from different viewpoints in loop closure [xiao2023semantic_loop_closure].
We have also made extensive explorations in the field of wide-FoV SLAM in the past. Building on the VINS series of works, we proposed LF-VIO [wang2022lf], the first visual-inertial tightly coupled framework that can adapt to radial field-of-view inputs exceeding 180 degrees. Later, based on LF-VIO, we expanded it into a complete SLAM framework, LF-VISLAM [wang2023lfvislam]. Similarly, building on LF-VIO, we proposed LF-PGVIO [wang2023lfvislam, wang2024lfpgvio], which introduced line features into the field of wide-field-of-view SLAM using geodesic constraints, achieving better positioning accuracy. In this paper, however, we focus on the impact of wide-field-of-view visual inputs on the construction of the SLAM problem itself, based on another equally mature SLAM framework, the monocular mode of ORB-SLAM3, and propose P2U-SLAM that can utilize point and pose uncertainties.
@@
Although there have been plenty of VIO and SLAM frameworks that support wide-FoV cameras, almost all of them pay no attention to the uncertainty of variables observed multiple times by wide-FoV cameras and the corresponding variables’ properties change during the location process. Some methods [qin2018vins, hu2019indoor, wang2022lf, matsuki2018omnidirectional] cannot update the uncertainty of variables continuously due to the limitation of direct methods or the optical flow tracking in the matching process. We point out that the methods based on descriptors have the potential to calculate and utilize the uncertainty of 3D points and 6D poses simultaneously. To make use of the potential of descriptors and overcome the limitations introduced by changes in variables’ properties, we propose the first visual wide-FoV SLAM system utilizing point- and pose uncertainty based on ORB features.
@@ III Camera Model
The camera model describes the measurement process of the 3D coordinate of a map point and outputs a projection pixel on the image plane as its measurement result. Many SLAM systems rectify the images to adapt to the pinhole camera model, while the way of distortion rectification cannot be flexibly transferred to a wide-FoV camera, whose FoV can even surpass 180180°. Equidistance projection and other non-perspective projection camera models are more suitable for calibration of wide-FoV cameras [kannala2006generic] since the rectification of wide-FoV images is unnecessary. Considering that the wide-FoV camera intrinsic parameters provided by different datasets may come from different camera models, we provide two popular camera models for wide-FoV cameras, which are the Taylor camera model [scaramuzza2006toolbox] and the Kannala-Brandt fisheye model [kannala2006generic].
@@ III-A Taylor Camera Model
As described in [hu2019indoor, scaramuzza2006toolbox], the projection function between a map point $\mathbf{p}$ with camera coordinate $(x,y,z)^{T}$ and its projection pixel $\mathbf{u}_{\pi}$ with coordinate $(u,v)^{T}$ can be written as
@@
$\rho(\theta)=a_{0}+a_{1}\cdot\theta^{1}+a_{2}\cdot\theta^{2}+a_{3}\cdot\theta^{3}+$.,
@@
[…]
@@
$z_{p}(r)=b_{0}+b_{1}\cdot r^{1}+b_{2}\cdot\ r^{2}+b_{3}\cdot r^{3}+$.,
@@
It should be noted that the Jacobian matrix of the camera projection function can be calculated via chain rules as Eq. (5), which matters in motion estimation and calculation of point uncertainty.